# LEGO-Puzzles: How Good Are MLLMs at Multi-Step Spatial Reasoning?

### v1

@@ 4.2 Main Results
We include evaluation results for spatial understanding, sequential reasoning, and generation in Table 1, Table 2, and Table 3. We summarize key findings as below.
Challenges of LEGO-Puzzles. Our findings indicate that human experts consistently achieve significantly higher overall performance (93.6%), as shown in Table 2. In contrast, current MLLMs fall short, with even the most advanced models, Gemini-2.0-Flash and GPT-4o, trailing over 30% behind human performance across all tasks. This persistent gap highlights the need for comprehensive and substantial improvements in our LEGO-Puzzles.
@@
Gap between Open-source and Proprietary Models. There is a significant gap between open-source and proprietary MLLMs in both spatial understanding and sequential reasoning abilities. Most open-source MLLMs perform only marginally better than Random, while leading proprietary models, such as Gemini-2.0-Flash and GPT-4o, exhibit strong spatial reasoning capabilities, achieving overall accuracies of 54.0% and 57.7%, respectively.
Model Performance in Different Tasks. In the Height task, where height relationships are complicated by the interplay between 2D and 3D perspectives, most models (11/20) perform worse than Random, with even human experts achieving significantly lower scores than in other tasks. In the Rotation and Rotation Status tasks, our findings indicate that models exhibit limited sensitivity to rotation-related recognition, achieving low scores, with 7 out of 20 models performing below Random in both tasks. Conversely, most models achieve accuracy above the critical threshold in the Multiview task, indicating that existing MLLMs possess basic spatial modeling capabilities. However, MLLMs’ performance is notably weaker in multi-step sequential reasoning tasks such as Ordering and Outlier compared to single-step spatial reasoning tasks like Dependency and Next-Step. This disparity highlights the models’ limitations in capturing long-range dependencies and executing effective sequential reasoning.
@@
In conclusion, our LEGO-Puzzles highlights both the spatial understanding and sequential reasoning abilities of MLLMs. Results in Table 1 demonstrate that GPT-4o achieves the highest performance. However, the overall results suggest significant room for improvement, particularly in domains involving relative relationships, rotation perception, and long-range sequential reasoning.
@@
App IF App IF App IF App IF App IF
Rotation* 2.30 1.65 0.95 0.80 2.10 0.00 0.00 0.00 0.10 0.00
Multiview* 1.80 1.35 2.25 0.45 2.10 0.00 0.00 0.00 0.05 0.00
Position* 3.00 1.40 3.00 1.10 0.65 0.00 0.00 0.00 0.00 0.00
Dependency* 1.85 1.25 0.55 0.25 0.65 0.00 0.00 0.00 0.00 0.00
Next-Step* 1.80 0.20 0.55 0.20 2.10 0.00 0.00 0.00 0.00 0.00
Overall 2.15 1.17 1.46 0.56 1.52 0.00 0.00 0.00 0.03 0.00
Table 3: Evaluation on Generation. We conduct human-based evaluation to assess the “Appearance” (App) and “Instruction Following” (IF) scores of Gemini-2.0-Flash, GPT-4o, Emu2, GILL, and Anole, using a scoring scale from 0 to 3 for both dimensions.
4.3 Image Generation Evaluation
As mentioned in Section 3.1, we evaluate image generation ability across several tasks related to spatial understanding (Rotation* and Multiview*) and single-step sequential reasoning (Position*, Dependency* and Next-Step*) as part of the Generation assessment. As shown in Table 3, all MLLMs struggle to simultaneously maintain appearance identity and strictly adhere to user instructions when generating image answers across all tasks. This poor performance indicates that existing MLLMs are ineffective at visualizing spatial understanding and sequential reasoning capabilities, underscoring the challenge of integrating multimodal information effectively.
@@ 4.4 Exploring Multi-Step Sequential Reasoning
Experiments in Section 4.2 show that current MLLMs perform poorly when extending single-step sequential reasoning QA to multi-step tasks. To further investigate the underlying reasons for these performance variations in sequential reasoning tasks, we design a fine-grained sequential reasoning task called Next-$k$-Step, which explicitly controls the number of steps required to complete the task.
Experimental Setup. Next-$k$-Step builds upon our single-step sequential reasoning task, Next-Step, and requires MLLMs to identify the correct LEGO object by sequentially adding $k$ additional LEGO pieces to the current LEGO object. We set $k=1,2,3,4,5$ and construct 20 test cases for each $k$ value. Specifically, we input current LEGO object ($x_{1}$), next $k$ LEGO pieces ($x_{2},x_{3},\dots,x_{k+1}$) and the target LEGO object ($x_{k+2}$), along with the corresponding text instructions into MLLMs, excepting them to generate the correct answer from four options (A, B, C, D). Additionally, to investigate the effectiveness of the widely adopted Chain of Thought (CoT) approach from the LLM community in enhancing multi-step sequential reasoning, we designed experiments comparing model performance under two conditions: standard prompting (without CoT) and explicit step-by-step reasoning (with CoT). We conduct experiments using the four top-performing models on the Next-Step task, GPT-4o, Gemini-2.0-Flash, Qwen-2.5-72B, and Internvl-2.5-78B.
Setting GPT-4o Gemini-2.0-Flash Qwen-2.5-72B Internvl-2.5-78B
@@
Performance Degradation when $k$ Increases. As shown in Table 4, the relationship between accuracy and the number of reasoning steps varies across models. GPT-4o and Gemini-2.0-Flash exhibit a clear performance decline as $k$ increases. These results align with the performance discrepancy between single-step and multi-step sequential reasoning in Table 1, further demonstrating that current MLLMs struggle to handle multi-step sequential relationships requiring iterative reasoning. A key challenge lies in the accumulation of errors as reasoning steps increase. Each intermediate inference introduces potential deviations that can compound over multiple steps, leading to significant inconsistencies in the final predictions. Additionally, MLLMs might lack an explicit visual memory mechanism compared to LLMs in language memory, making it difficult to coherently track and integrate positional changes throughout the reasoning process. Surprisingly, we observe that Qwen-2.5-72B achieves stable and relatively consistent accuracy scores (around 65.0%) across all values of $k$, even when $k$ is increased to seven. And for Internvl-2.5-78B, accuracy scores are close to random guessing.
Limited Effectiveness of Chain-of-Thought (CoT). By applying CoT prompting, we observe significant improvements when $k=1$ for GPT-4o and InternVL-2.5-78B. However, this effect diminishes for $k\geq 2$, where accuracy even declines dramatically. This is because these MLLMs perform worse than random guessing (25%) when dealing with longer $k$ steps, with accuracies of 5% for GPT-4o and 10% for InternVL-2.5-78B. For other MLLMs like Gemini-2.0-Flash and Qwen-2.5-72B, CoT prompting does not provide obvious benefits, as they fail to perform genuine step-by-step reasoning in their CoT responses.
@@
Table 5: Pearson Correlation Coefficients (PCC) and P-values for Height and Adjacency Tasks
4.5 Consistency Compared with Natural Dataset
Besides its high scalability as a virtual framework, LEGO-Puzzles also demonstrates strong consistency with the natural environment. To verify this, we compare LEGO-Puzzles with 3DSRBench [38], which includes several similar tasks (Height in LEGO-Puzzles and 3DSRBench, Adjacency in LEGO-Puzzles and Location in 3DSRBench) but focuses on real-world domain images. Specifically, we evaluate all proprietary models tested on LEGO-Puzzles within the 3DSRBench dataset and compute the Pearson correlation coefficient [10] to measure the accuracy correlation between these two benchmarks. The results in Table 5 indicate that model performance on LEGO-Puzzles reliably reflects trends observed in natural data.
Figure 5: Task Similarity Heatmap. The heatmap illustrates the pairwise correlation between tasks in our benchmark, measured using SRCC, PLCC, and R² scores.
4.6 Task Similarity
In this subsection, we analyze task similarity in our benchmark by calculating the average rank correlation between each task and all others, as proposed by Zhang et al. [61]. The similarity score for each task is derived from the average correlation between its ranking and those of all other tasks, using three metrics: Spearman Rank Correlation Coefficient (SRCC), Pearson Linear Correlation Coefficient (PLCC), and R² Score (R-squared Coefficient of Determination).
The results in Figure 5 show that only a few task pairs have strong correlations, such as Next Step to Dependency, Multi-step to Single-step Sequential Reasoning, and Spatial Understanding. This is because they either share similar image inputs or integrate step-wise logical progression and spatial comprehension. However, most tasks exhibit moderate to low correlations, ensuring benchmark diversity. More than half of the task pairs have an SRCC between 0.3 and 0.6, indicating limited dependency among tasks.
Overall, our benchmark offers a balanced assessment of MLLMs across various reasoning skills. While some tasks show strong correlations due to conceptual overlap, the majority remain sufficiently independent, providing a comprehensive and distinctive evaluation framework.
5 Error Analysis
In this section, we conduct a detailed error analysis during the evaluation process of our benchmark, providing insights into model behaviors and shortcomings.
Figure 6: Visualization of sample failure cases in Height and Ordering. The Ground Truth answer is marked in blue, while the MLLM’s answer is marked in red. Note: The questions above are slightly simplified for clarity and brevity.
Failure in Ordering Task. As shown in Table 1, several open-source models (4/14) completely fail in the Ordering task, with scoring zero. Ordering task requires MLLMs to enable multi-step reasoning ability. Despite explicitly specifying the required answer format in the prompt (e.g., a sequence such as “BACD”), some models were unable to generate a valid response and instead produced arbitrary outputs. For instance, several models (InternVL2.5-8B, Emu3, MiniCPM-V-2-6, and LLava-OneVision-7B) exhibited strong biases, frequently defaulting to a single letter rather than a complete sequence. In extreme cases, models provided nearly identical responses across all test instances, such as Emu3, which answered “B” for 98 out of 100 test cases, demonstrating a lack of genuine reasoning ability. These results indicate that many open-source models struggle with sequence generation and constrained output formatting, suggesting potential issues in their ability to follow structured prompts for reasoning tasks. These biases in responses further suggest that models may be overly reliant on spurious correlations in training data rather than understanding the stepwise dependencies of a given sequence.
Challenges in Height Perception: 2D vs. 3D Understanding. In the Height task, we observe that most models (11/20) achieve scores lower than Random. We provide some failure cases in Figure 6, which exhibit noticeable 2D and 3D optical illusions. Since MLLMs are primarily trained on images with a predominantly 2D viewpoint, the discrepancy between 2D and 3D spatial understanding in our Height task often causes MLLMs to answer questions based on a 2D projection rather than a true 3D perspective. Even when we construct instruction prompts that explicitly require MLLMs to comprehend 3D spatial relationships, GPT-4o, the top-performing model, still fails to achieve human-level performance. This observation highlights the tendency of MLLMs to rely on 2D spatial priors during inference, suggesting the need for further research in 3D understanding training.
Figure 7: Qualitative image generation results for Rotation* and Multiview* tasks. Note: The questions above are slightly simplified for clarity and brevity.
Weak Appearance Consistency and Instruction Following in Image Generation. Our evaluation of MLLM-generated images reveals substantial differences in instruction following and reasoning-based image synthesis when processing sequential visual inputs. As shown in Table 7, open-source models struggle significantly in both appearance consistency (App) and instruction-following (IF), while proprietary models demonstrate varying degrees of success. Among the proprietary models, Gemini-2.0-Flash exhibits the strongest performance in both appearance and instruction adherence. It effectively follows input constraints and maintains high appearance fidelity, often editing the given image rather than generating a completely new one. This suggests that Gemini-2.0-Flash has a stronger spatial consistency mechanism, enabling it to make precise modifications while preserving structural coherence. For GPT-4o, the results suggests that GPT-4o may not directly edit the input image but instead interpret its semantic content and generate a new image based on textual understanding. The differences in appearance fidelity and instruction following indicate that GPT-4o’s generation process might involve reconstructing the scene conceptually rather than modifying the original image step by step. While this allows it to maintain conceptual relevance, its output often deviates in style and structure from the original input, leading to lower appearance fidelity compared to Gemini-2.0-Flash. Moreover, its instruction following ability remains inconsistent, particularly for tasks requiring fine-grained reasoning. For open-source models, Emu2 exhibits some capability in preserving visual appearance but fails entirely in instruction-following, treating the task as mere image reconstruction rather than reasoning-based generation. It struggles with spatial dependencies and sequential modifications, making it ineffective for reasoning-intensive tasks. GILL and Anole perform the worst, failing to generate relevant outputs in nearly all cases. Their instruction-following scores are close to zero, and their generated images are often completely unrelated to the expected result. This highlights a fundamental limitation in their ability to process sequential visual transformations, making them unsuitable for complex, instruction-driven image generation. We provide some failure cases in Figure 7. These findings emphasize the fundamental challenges in spatial and sequential reasoning within open-source MLLMs. While Gemini-2.0-Flash shows the most precise adherence to instructions and image editing capabilities, GPT-4o tends to generate semantically relevant but visually divergent outputs. Open-source models, by contrast, still lack robust mechanisms for sequential reasoning, underscoring the need for advancements in instruction-following and reasoning-aware image generation techniques.

### v2

@@ 4.2 Main Results
We include evaluation results for spatial understanding and sequential reasoning in Tab. 1 and Tab. 2. We summarize key findings as below.
@@
Gap between Open-source and Proprietary Models. There is a significant gap between open-source and proprietary MLLMs in both spatial understanding and sequential reasoning abilities as shown in Tab. 1. Most open-source MLLMs perform only marginally better than Random, while leading proprietary models, such as Gemini-2.0-Flash and GPT-4o, exhibit strong spatial reasoning capabilities, achieving overall accuracies of 54.0% and 57.7%, respectively.
Gap between Human and MLLMs. To study the performance gap between human and MLLMs on LEGO-Puzzles tasks, we randomly select 20 questions from each task to create LEGO-Puzzles-Lite, resulting in a total of 220 QA pairs, and use this subset for investigation. We invited 30 human experts to solve these questions, ensuring a diverse and representative evaluation of human-level reasoning. Our findings indicate that human experts consistently achieve significantly higher overall performance (93.6%), as shown in Tab. 2. In contrast, current MLLMs fall short, with even the most advanced models, Gemini-2.0-Flash and GPT-4o, trailing over 30% behind human performance across all tasks. This persistent gap highlights the need for comprehensive and substantial improvements in our LEGO-Puzzles.
Challenges of LEGO-Puzzles. 1). Spatial Understanding. Our results reveal that MLLMs struggle significantly with tasks requiring true 3D spatial reasoning. In the Height task, we deliberately construct cases where interpreting the image from a 2D viewpoint leads to a different answer than interpreting it from a true 3D perspective. As shown in Tab. 1, most models (11/20) perform worse than Random. These results suggest that MLLMs tend to defaultly answer questions based on a 2D projection rather than a true 3D perspective. This observation highlights their reliance on 2D spatial priors during inference, underscoring the need for further research on equipping models with robust 3D spatial reasoning capabilities. Similarly, in the Rotation task, which requires identifying the rotation angle of an object, 7 out of 20 models fall below Random, with most failing to reach even 40% accuracy. This further indicates that current MLLMs struggle to perceive and distinguish object orientation changes reliably. 2). Sequential Reasoning. LEGO-Puzzles also reveals substantial limitations in MLLMs’ sequential reasoning capabilities, particularly when multiple reasoning steps are required. While performance on single-step tasks such as Dependency and Next-Step is relatively stronger, models struggle considerably with multi-step tasks. As shown in Tab. 1, almost half of the models score below $\uparrow$ Random in the Ordering task, with some models (e.g., InternVL2.5-8B, LLaVA-OneVision-7B, EMU3) failing completely. Similar trends are observed in the Backwards task, where 9 out of 14 open-source models perform below $\uparrow$ Random. To further understand the effect of step length, we conduct the Next-k-Step experiment in Sec. 4.4 which systematically evaluates model accuracy under increasing reasoning steps.
@@
In conclusion, our LEGO-Puzzles highlights both the spatial understanding and sequential reasoning abilities of MLLMs. The overall results suggest significant room for improvement, particularly in domains involving relative relationships, rotation perception, and long-range sequential reasoning.
4.3 Image Generation Evaluation
As mentioned in Sec. 1, to further investigate whether MLLMs can transfer their spatial understanding and reasoning abilities to image generation, we construct a set of generation tasks based on LEGO-Puzzles. Specifically, we convert the original multiple-choice format in LEGO-Puzzles into image generation tasks, where models are required to directly produce a visual output rather than selecting from given options. We select five tasks—Rotation* and Multiview* from spatial understanding, and Position*, Dependency*, and Next-Step* from single-step sequential reasoning—resulting in a total of 100 questions.
We evaluate the open-source models Emu2 [46], GILL [22], and Anole [9], as well as the proprietary models GPT-4o\$$\triangle$\$$\triangle$\$footnotetext: refers to the version of GPT-4o released prior to March 6, 2025., GPT-4o and Gemini-2.0-Flash, all of which support long-range sequence input and image output. For evaluation, traditional metrics such as FID [19], CLIPScore [18, 14], and X-IQE [6] mainly assess image fidelity or cross-modal alignment, often relying on pre-trained model priors or fixed scoring heuristics. They struggle to capture the fine-grained spatial accuracy required in LEGO assembly tasks, where even small errors—such as misaligned parts or incorrect orientations—can invalidate the result. Furthermore, many recent multimodal evaluation metrics depend on GPT-based models [34], introducing uncontrollable bias into the evaluation process. Therefore, we enlist 5 human experts to assess model performance across two dimensions: appearance similarity and instruction following. Each aspect is rated on a scale from 0 to 3.
Table 3: Evaluation on generation. We conduct human-based evaluation to assess the “Appearance” (App) and “Instruction Following” (IF) scores of GPT-4o, Gemini-2.0-Flash, GPT-4o, Emu2, GILL, and Anole, using a scoring scale from 0 to 3 for both dimensions.
@@ 4.3 Image Generation Evaluation
App IF App IF App IF App IF App IF App IF
Rotation* 2.35 1.75 2.05 1.45 0.65 0.50 1.70 0.00 0.00 0.00 0.10 0.00
Multiview* 2.30 2.00 2.40 1.65 1.95 0.40 1.65 0.25 0.00 0.00 0.05 0.00
Position* 2.30 1.65 2.85 1.30 2.95 1.00 0.50 0.00 0.00 0.00 0.00 0.00
Dependency* 2.05 2.00 1.70 0.95 0.35 0.05 0.55 0.00 0.00 0.00 0.00 0.00
Next-Step* 2.25 1.45 1.75 0.05 0.30 0.00 0.05 0.00 0.00 0.00 0.00 0.00
Overall 2.25 1.77 2.15 1.08 1.24 0.39 0.89 0.05 0.00 0.00 0.03 0.00
Tab. 3 presents the human evaluation results across five generation tasks. Overall, proprietary models outperform open-source ones in both appearance consistency (App) and instruction adherence (IF). Among them, GPT-4o achieves the highest overall scores (App: 2.25, IF: 1.77), followed by Gemini-2.0-Flash (App: 2.15, IF: 1.08). However, both models show clear room for improvement in instruction following, scoring only 1.77 and 1.08 out of 3, respectively. For GPT-4o, the results suggest that the model may not directly edit the input image, but instead reinterpret the scene semantically and regenerate it based on textual understanding. This leads to lower appearance consistency (App: 1.24), reflecting a conceptual reconstruction process rather than precise visual editing. Among open-source models, Emu2 shows some ability to preserve visual appearance (App: 0.89) but fails almost entirely in instruction following (IF: 0.05), treating the task more as image replication than reasoning-based generation. GILL and Anole perform the worst, with near-zero scores across all tasks and frequently irrelevant outputs. The qualitative results are shown in Fig. 3.
Figure 3: Qualitative visual results for image generation tasks. Note: The questions above are slightly simplified for clarity and brevity.
Overall, these results show that current models—especially open-source ones—struggle significantly with instruction-grounded image generation, highlighting the challenges of spatially grounded visual synthesis.
@@ 4.4 Exploring Multi-Step Sequential Reasoning
Experiments in Sec. 4.2 show that MLLMs perform noticeably worse on multi-step sequential reasoning tasks compared to single-step ones. To systematically investigate how reasoning performance is affected by the number of steps involved, we design a fine-grained task called Next-$k$-Step. This task extends the original Next-Step setting in LEGO-Puzzles by explicitly controlling the number of steps $k$ required to reach the target assembly state.
Experimental Setup. In contrast to Next-Step in LEGO-Puzzles, Next-$k$-Step requires MLLMs to identify the correct LEGO object after sequentially adding $k$ additional LEGO pieces to the current LEGO object. We set $k=1,2,3,4,5$ and construct 20 test cases for each $k$ value. Specifically, each input includes the current LEGO object ($x_{1}$), the next $k$ LEGO pieces ($x_{2},x_{3},\dots,x_{k+1}$) and the target LEGO object ($x_{k+2}$), along with the corresponding text instructions. The model is expected to select the correct answer from four options (A, B, C, D). To further explore whether the widely used Chain-of-Thought (CoT) prompting strategy can enhance performance in multi-step reasoning, we evaluate model accuracy under two prompting conditions: standard prompting (without CoT) and explicit step-by-step reasoning (with CoT). We conduct this experiment using the four top-performing models from the original Next-Step task: GPT-4o, Gemini-2.0-Flash, Qwen2.5-VL-72B, and InternVL2.5-78B.
Performance Degradation when $k$ Increases. As shown in Tab. 4, GPT-4o and Gemini-2.0-Flash exhibit clear performance degradation as the number of reasoning steps $k$ increases, reflecting their difficulty in handling multi-step sequential reasoning. In contrast, Qwen2.5-VL-72B maintains stable accuracy ( 65%) across all $k$ values, suggesting stronger robustness to multi-step reasoning. InternVL2.5-78B performs near the random baseline (25%) regardless of $k$, indicating limited overall effectiveness. These results suggest that most current MLLMs lack the capacity to track and integrate spatial transformations over multiple steps, with accumulated errors leading to inconsistent predictions in longer reasoning chains.
Setting GPT-4o Gemini-2.0-Flash Qwen2.5-VL-72B InternVL2.5-78B
@@
Limited Effectiveness of Chain-of-Thought (CoT). As shown in Tab. 4, CoT provides clear gains at $k=1$ for GPT-4o ($\uparrow$30%) and InternVL2.5-78B ($\uparrow$20%), and a modest improvement for GPT-4o at $k=2$ ($\uparrow$10%). However, for $k\geq 2$, its effect vanishes or worsens—GPT-4o drops to 0% by $k=4$, and InternVL2.5-78B stays near or below the random baseline (25%). For Gemini-2.0-Flash and Qwen2.5-VL-72B, CoT yields no consistent gain. In particular, Qwen2.5-VL-72B performs stably with or without CoT, suggesting robustness without reliance on explicit step-by-step prompting. Overall, CoT offers limited help in multi-step spatial reasoning, especially as task complexity increases.
5 Discussion
While LEGO-Puzzles is built on rendered data, it aims to evaluate fundamental spatial reasoning capabilities that are also essential in real-world scenarios. To assess its generalizability beyond synthetic environments, we compare model performance on LEGO-Puzzles with 3DSRBench [37], a benchmark based on natural images. Both datasets contain conceptually similar tasks—specifically, the Height task in LEGO-Puzzles aligns with Height in 3DSRBench, and Adjacency in LEGO-Puzzles corresponds to the Location task in 3DSRBench.
Table 5: Pearson correlation coefficients (PCC) and p-values for height and adjacency tasks.
@@ 5 Discussion
We evaluate all proprietary models tested in LEGO-Puzzles on the corresponding tasks in 3DSRBench and compute the Pearson correlation coefficient [10] to measure consistency in performance across the two datasets. As shown in Tab. 5, the results reveal strong positive correlations: 0.93 for Height and 0.98 for Adjacency, both statistically significant ($p<0.01$).
These findings suggest that LEGO-Puzzles not only offers high scalability and precise control in synthetic settings but also captures spatial reasoning patterns that generalize well to natural images. This validates its utility as a proxy for evaluating real-world spatial understanding.